# Contribute to balance, wire in accordance: Emergence of backpropagation from a simple, bio-plausible neuroplasticity rule

### v1

@@ 2.2 Neural firing
We utilize an all-to-all connected neural network to establish our model (Figure 1A), which provides a high degree of generality. The firing activities of neurons are represented as an $n$-dimensional vector $\mathbf{x}(t)$, where the $i^{\text{th}}$ element $x_{i}(t)$ reflects the firing activity of the $i^{\text{th}}$ neuron at time step $t$. The firing activity depends on the nonlinear activation function $\sigma$ and the inputs to each neuron $\mathbf{s}(t)=\mathbf{Wx}(t)$, where $\mathbf{W}$ is the connectivity matrix. The element $w_{ij}$ in $\mathbf{W}$ indicates the strength of the synapse received by neuron $i$ from neuron $j$. Consequently, the network dynamics and steady state (fixed points) can be described as follows:
$\mathbf{x}(t+1)=\sigma(\mathbf{W}\mathbf{x}(t)),\quad\mathbf{x}=\sigma(\mathbf{W}\mathbf{x})$ (1)
@@ 2.3 Neural credits
In addition to the firing rate, we hypothesize that each neuron encodes a second quantity that we refer to as credit. Credits of neurons are represented as an $n$-dimensional vector $\mathbf{c}(t)$, where the $i^{\text{th}}$ element $c_{i}(t)$ reflects the amount of credit possessed by the $i^{\text{th}}$ neuron at time step $t$.
@@
The dynamics of credit redistribution for each presynaptic neuron $j$ can be described as $c_{j}(t+1)=\sum_{i}g_{ji}(t)$, with the summation occuring over all of its postsynaptic neurons $i$. Consequently, the network dynamics of credit redistribution and the corresponding fixed points can be described as:
$\mathbf{c}(t+1)=\mathbf{J}\mathbf{c}(t)=\begin{bmatrix}f(s_{1})w_{11}&f(s_{2})w_{21}&\cdots&f(s_{n})w_{n1}\\ f(s_{1})w_{12}&f(s_{2})w_{22}&\cdots&f(s_{n})w_{n2}\\ \vdots&\vdots&\ddots&\vdots\\ f(s_{1})w_{1n}&f(s_{2})w_{2n}&\cdots&f(s_{n})w_{nn}\end{bmatrix}\begin{bmatrix}c_{1}(t)\\ c_{2}(t)\\ \vdots\\ c_{n}(t)\end{bmatrix},\quad\mathbf{c}=\mathbf{J}\mathbf{c}$ (3)
The matrix $\mathbf{J}$ summarizes these dynamics. Here, $t$ is the time index specifically for the dynamics of credit redistribution, which is independent from the time index in neural firing dynamics, though the same notation is used for simplicity. The time for credit redistribution is expected to be in the range of a second or a few seconds, based on the known biological features of retrograde signaling.
@@ 2.4 Neuroplasticity rule
$\Delta w_{ij}=c_{i}f(s_{i})x_{j},\quad\text{i.e.}\quad\Delta\mathbf{W}=(\mathbf{c}\odot f(\mathbf{s}))\mathbf{x}^{T}$ (4)
@@ 2.5 Timescales of the model
Our model operates on three sequentially slower timescales: firing activity on the millisecond scale, credit redistribution on the seconds scale, and neural plasticity from seconds to minutes. This is depicted graphically in the raster plot of figure 2A, where each vertical line represents an update event rather than a spike. When the system undergoes significant changes (for instance, owing to the arrival of a new input), events are densely distributed. As the system stabilizes and converges, the events gradually disappear. Due to the hierarchical nature of these different timescales, we posit that the time it takes for one level to converge is much shorter than the update time interval of the subsequent level. Consequently, most credit update events occur when network firing has already converged, and most weight update events occur when credit redistribution has converged. In other words, equation 3 is based on the equilibrium state of equation 1, while equation 4 is based on the equilibrium state of equation 3.
@@
The time windows for input clamping (blue shadow) and output clamping (yellow shadow) are illustrated in figure 2A. An overlap (green shadow) between these two time windows is necessary because, outside this overlap, either $\mathbf{x}$ or $\mathbf{c}$ will quickly converge to zero. Additionally, the overlap must be long enough for neuroplasticity to occur. Therefore, the durations of both input and output clamping should align with the neuroplasticity timescale, which is from seconds to longer. As a result of these multiple timescales, the system automatically distinguishes between the prediction phase (only input clamping) and the learning phase (overlapped input and output clamping), thereby avoiding the need to explicitly introduce distinct phases. From a biological perspective, we hypothesize that these multiple hierarchical timescales are orchestrated by neural oscillations (see also section 5).
Figure 2: Three timescales of network operation. (A) Each row represents updates to the corresponding quantity; timescales are progressively slower from top to bottom. Each vertical line represents an update event; densely distributed events correspond to rapid updates; and update events disappear as quantities converge to steady-state values. Learning occurs with input clamping on neural firing (top row, blue shading) and output clamping on credit values (middle row, yellow shading), as long as there is a sufficiently long period of overlap (bottom row, green shading). (B,C) Input and output clamping in an all-to-all connected ANN, and a layered ANN, respectively.
@@ 2.6 Summary of the model
In summary, our new framework operates following three update rules: (1) Neural firing dynamics at the fastest timescale: $\mathbf{x}(t+1)=\sigma(\mathbf{W}\mathbf{x}(t))$; (2) Credit redistribution dynamics at the intermediate timescale: $\mathbf{c}(t+1)=\mathbf{J}\mathbf{c}(t)$; (3) Neural plasticity dynamics at the slowest timescale: $\mathbf{W}(t+1)=(\mathbf{c}\odot f(\mathbf{s}))\mathbf{x}^{T}+\mathbf{W}(t)$
@@ Acknowledgements
This work was supported by a JHU Discovery Award co-funded with the One Neuro Initiative (SPM). We thank Hokin Deng (JHU) for making helpful suggestions about the title of the manuscript.

### v2

@@ 2.2 Neural firing
We utilize an all-to-all connected neural network to establish our model (Figure 1A), which provides a high degree of generality. The firing activities of neurons are represented as an $n$-dimensional vector $\mathbf{x}(t^{\prime\prime})$, where the $i^{\text{th}}$ element $x_{i}(t^{\prime\prime})$ reflects the firing activity of the $i^{\text{th}}$ neuron at time step $t^{\prime\prime}$. The firing activity depends on the nonlinear activation function $\sigma$ and the inputs to each neuron $\mathbf{s}(t^{\prime\prime})=\mathbf{Wx}(t^{\prime\prime})$, where $\mathbf{W}$ is the connectivity matrix. The element $w_{ij}$ in $\mathbf{W}$ indicates the strength of the synapse received by neuron $i$ from neuron $j$. Consequently, the network dynamics and steady state (fixed points) can be described as follows:
$\mathbf{x}(t^{\prime\prime}+1)=\sigma(\mathbf{W}\mathbf{x}(t^{\prime\prime})),\quad\mathbf{x}=\sigma(\mathbf{W}\mathbf{x})$ (1)
@@ 2.3 Neural credits
In addition to the firing rate, we hypothesize that each neuron encodes a second quantity that we refer to as credit. Credits of neurons are represented as an $n$-dimensional vector $\mathbf{c}(t^{\prime})$, where the $i^{\text{th}}$ element $c_{i}(t^{\prime})$ reflects the amount of credit possessed by the $i^{\text{th}}$ neuron at time step $t^{\prime}$.
@@
The dynamics of credit redistribution for each presynaptic neuron $j$ can be described as $c_{j}(t^{\prime}+1)=\sum_{i}g_{ji}(t^{\prime})$, with the summation occuring over all of its postsynaptic neurons $i$. Consequently, the network dynamics of credit redistribution and the corresponding fixed points can be described as:
$\mathbf{c}(t^{\prime}+1)=\mathbf{J}\mathbf{c}(t^{\prime})=\begin{bmatrix}f(s_{1})w_{11}&f(s_{2})w_{21}&\cdots&f(s_{n})w_{n1}\\ f(s_{1})w_{12}&f(s_{2})w_{22}&\cdots&f(s_{n})w_{n2}\\ \vdots&\vdots&\ddots&\vdots\\ f(s_{1})w_{1n}&f(s_{2})w_{2n}&\cdots&f(s_{n})w_{nn}\end{bmatrix}\begin{bmatrix}c_{1}(t^{\prime})\\ c_{2}(t^{\prime})\\ \vdots\\ c_{n}(t^{\prime})\end{bmatrix},\quad\mathbf{c}=\mathbf{J}\mathbf{c}$ (3)
The matrix $\mathbf{J}$ summarizes these dynamics. Here, $t^{\prime}$ is the time index specifically for the dynamics of credit redistribution, which is independent from the time index in neural firing dynamics $t^{\prime\prime}$. The time for credit redistribution is expected to be in the range of a second or a few seconds, based on the known biological features of retrograde signaling.
@@ 2.4 Neuroplasticity rule
$\Delta w_{ij}=c_{i}f(s_{i})x_{j},\quad\text{i.e.}\quad\mathbf{W}(t+1)=(\mathbf{c}\odot f(\mathbf{s}))\mathbf{x}^{T}+\mathbf{W}(t)$ (4)
@@ 2.5 Timescales of the model
Our model operates on three sequentially slower timescales: firing activity on the millisecond scale, credit redistribution on the seconds scale, and neural plasticity from seconds to minutes. This is depicted graphically in the raster plot of figure 2A, where each update mechanism is shown at its respective timescale using distinct time bin sizes for reference. When the system undergoes significant changes (for instance, owing to the arrival of a new input), the amplitude of curve is high. As the system stabilizes and converges, the amplitude gradually decreases to zero. Due to the hierarchical nature of these different timescales, we posit that the time it takes for one level to converge is much shorter than the update time interval of the subsequent level. Consequently, most changes due to credit update (yellow shadow under the curve) occur when network firing has already converged, and most weight changes (green shadow) occur when credit redistribution has converged. In other words, equation 3 is based on the equilibrium state of equation 1, while equation 4 is based on the equilibrium state of equation 3.
@@
The time windows for input clamping (signal 1) and output clamping (signal 2) are illustrated in figure 2A. An overlap between these two time windows is necessary because, outside this overlap, either $\mathbf{x}$ or $\mathbf{c}$ will quickly converge to zero. Additionally, the overlap must be long enough for neuroplasticity to occur. Therefore, the durations of both input and output clamping should align with the neuroplasticity timescale, which is from seconds to longer. As a result of these multiple timescales, the system automatically distinguishes between the prediction phase (only input clamping) and the learning phase (overlapped input and output clamping), thereby avoiding the need to explicitly introduce distinct phases. From a biological perspective, we hypothesize that these multiple hierarchical timescales are orchestrated by neural oscillations (see also section 5).
Figure 2: Three timescales of network operation. (A) Each row represents updates to the corresponding quantity; timescales are progressively faster from top to bottom. For the top three rows, each vertical line represents one timebin, and the curve represents the amplitude of the rate of updates. The area under the curve thus represents the total amount of change made to different quantities. As the quantity converges to steady-state, the corresponding curve value decreases to zero. Learning occurs with input clamping on neural firing (signal 1, blue) and output clamping on credit values (signal 2, yellow), as long as there is a sufficiently long period of overlap. (B,C) Input and output clamping in an all-to-all connected ANN, and a layered ANN, respectively.
@@ 2.6 Summary of the model
In summary, our new framework operates following three update rules: (1) Neural firing dynamics at the fastest timescale: $\mathbf{x}(t^{\prime\prime}+1)=\sigma(\mathbf{W}\mathbf{x}(t^{\prime\prime}))$; (2) Credit redistribution dynamics at the intermediate timescale: $\mathbf{c}(t^{\prime}+1)=\mathbf{J}\mathbf{c}(t^{\prime})$; (3) Neural plasticity dynamics at the slowest timescale: $\mathbf{W}(t+1)=(\mathbf{c}\odot f(\mathbf{s}))\mathbf{x}^{T}+\mathbf{W}(t)$
@@ Acknowledgements
This work was supported by funding from a JHU Discovery Award co-funded with the One Neuro Initiative (SPM), and from NIH grant 2R01EY027718 (SPM). We thank Hokin Deng (JHU) for making helpful suggestions about the title of the manuscript.